OpenStudy (anonymous):

the pet shelter feeds puppies twice a day and grown dogs once a day they have 38 puppies and 24 full grown dogs how many meals does the shelter feed everyday

OpenStudy (anonymous):

the f*ck are they pampering puppies so much for?

OpenStudy (anonymous):

XD i dont know

OpenStudy (anonymous):

Puppies x 2 + Dogs x 1. (38*2)+(24*1)=76 + 24 = 100

OpenStudy (anonymous):

what is this again *?